A druggist has two mixtures of a certain chemical and water, one containing 10% of the chemical and the other containing 40% of the chemical. How much of each mixture should he use to make 10 ounces that contain 25% of the chemical?
___oz of each
I'd appreciate any help. Thanks in advance!

Answer by checkley71(8403)

.10X+.40(10-X)=.25*10
.10X+4-.4X=2.5
.3X=2.5-4
.3X=1.5
X=1.5/.3
X=5 OZ. OF EACH MIXTURE IS NEEDED.
PROOF:
.10*5+.40*5=2.5
.5+2=2.5
2.5=2.5